You may need to use the appropriate technology to answer this question.

Consider the following hypothesis test.

H0: σ12 = σ22

Ha: σ12 ≠ σ22

(a)

What is your conclusion if

n1 = 21,

s12 = 2.2,

n2 = 26,

and

s22 = 1.0?

Use

α = 0.05

and the p-value approach.

Find the value of the test statistic.

Find the p-value. (Round your answer to four decimal places.)

p-value =

State your conclusion.

Reject H0. We cannot conclude that σ12 ≠ σ22.Do not reject H0. We cannot conclude that σ12 ≠ σ22.    Reject H0. We can conclude that σ12 ≠ σ22.Do not reject H0. We can conclude that σ12 ≠ σ22.

(b)

Repeat the test using the critical value approach.

Find the value of the test statistic.

State the critical values for the rejection rule. (Round your answers to two decimal places. If you are only using one tail, enter NONE for the unused tail.)

test statistic≤test statistic≥

State your conclusion.

Reject H0. We cannot conclude that σ12 ≠ σ22.Do not reject H0. We cannot conclude that σ12 ≠ σ22.    Reject H0. We can conclude that σ12 ≠ σ22.Do not reject H0. We can conclude that σ12 ≠ σ22.
Answer #1

Ans:

a)

Test statistic:

F=2.2/1.0=2.2

df1=21-1=20

df2=26-1=25

p-value(two tailed)=2*FDIST(2.2,20,25)=0.0633

As,p-value>0.05,so

Do not reject H0. We cannot conclude that σ12 ≠ σ22.

b)alpha/2=0.025

critical F values:

FINV(0.025,20,25)=2.30

FINV(0.975,20,25)=0.42

Reject H0,if F<=0.42 or F>=2.30

Do not reject H0. We cannot conclude that σ12 ≠ σ22.